Object X

 of mass M

 travels toward object Y

 of mass 2M

 in such a way that they collide. The table contains data about the velocities of object X

 and object Y

 immediately before the collision and immediately after the collision. What are the change in momentum of the two-object system and the change in momentum of each object from immediately before the collision to immediately after the collision?


1
Expert's answer
2021-10-06T08:15:33-0400

Change in momentum of the system


"m_1v_1 + m_2v_2 – m_1u_1 – m_2u_2 \\

M(\frac{2v_0}{3}-4v_0) + 2M(\frac{2v_0}{3}-(-v_0)) \\

\frac{-10v_0M}{3}+{10v_0M}{3}= 0"


Change m momentum of "x = M(\frac{2v_0}{3}-4v_0)= - \frac{10Mv_0}{3}"

Change m momentum of "y = 2M(\frac{2v_0}{3}-(-v_0))= \frac{+10Mv_0}{3}"


So, option A is correct